OpenStudy (anonymous):

What is the slope of this line? x=4

OpenStudy (anonymous):

Vertical!!!!

OpenStudy (anonymous):

Well it is vertical cx But what would ur x int. be Madeline? Quite easy

OpenStudy (anonymous):

idk i have a c in algebra XD idk what the heck im doing :P

OpenStudy (anonymous):

x=4 means ur x int is 4

OpenStudy (anonymous):

um. ok XD

OpenStudy (anonymous):

Basically if u just have the u draw a straight line up and down on that point on the line of the x axiz. Make sense?

OpenStudy (anonymous):

i guess haha XD thanks

OpenStudy (anonymous):

c:

OpenStudy (anonymous):

Woah i just realized i missed " u just have the x,"

OpenStudy (anonymous):



OpenStudy (anonymous):

haha doesnt matter what u missed cuz i still dunno what im doing XD LOL

OpenStudy (anonymous):

oh lord

OpenStudy (anonymous):

lol XD

OpenStudy (anonymous):

sorry for being so stupid :P XD

OpenStudy (anonymous):

just do rise over run

OpenStudy (accessdenied):

There are a few ways to think of slope: the rise of the graph per the run, or change in y per change in x. We figured this was a vertical line, so it can rise indefinitely without a change in x. Or perhaps using an equation of slope: m = (y2 - y1)/(x2 - x1) We take any two points on that line, say (4, 4) and (4, 0) (both have x -value 4 so they are on the line): m = (4 - 0)/(4 - 4) = 4/0 If this were 0/4, then we would have 0 slope. This is the typical situation when we have y defined rather than x. Division by zero, however, is not possible. We call this undefined.

OpenStudy (anonymous):

thank you!

OpenStudy (accessdenied):

You're welcome! :)